# Ball Lightning as a profound manifestation of the Dark Matter physics

### v1

@@ Abstract
Ball lighting (BL) has been observed for centuries. There are large number of books, review articles, and original scientific papers devoted to different aspects of BL phenomenon. Yet, the basic features of this phenomenon have never been explained by known physics. The main problem is the source which could power the dynamics of the BL. We advocate an idea that the dark matter in form of the axion quark nuggets (AQN) made of standard model quarks and gluons (similar to the old idea of the Witten’s strangelets) could internally generate the required power. The corresponding macroscopically large object in form of the AQN behaves as chameleon: it does not interact with the surrounding material in dilute environment and serves as perfect cold DM candidate. However, AQN becomes strongly interacting object in sufficiently dense environment. The AQN model was invented long ago without any relation to the BL physics. It was invented with a single motivation to explain the observed similarity $\Omega_{\rm DM}\sim\Omega_{\rm visible}$ between visible and DM components. This relation represents a very generic feature of this framework, not sensitive to any parameters of the construction. However, with the same set of parameters being fixed long ago this model is capable to address the key elements of the BL phenomenology, including the source of the energy powering the BL events. In particular, we argue that the visible size of BL, its typical life time, the frequency of appearance , etc are all consistent with suggested proposal when BL represents a profound manifestation of the DM physics represented by the AQN objects. We also formulate a unique possible test which can refute or unambiguously substantiate this unorthodox proposal on nature of BL.
@@ 1 Introduction
The title of this work seemingly includes two contradicting terms: the first one is “ball lighting”, which is very bright luminous object, though of unknown nature. The second term of the title is “dark matter” (DM) which is, by definition, must decouple from the radiation, as it cannot emit light. Indeed, the BL phenomenon has been known for centuries, see review papers [1, 2, 3] with large number of references therein. See also recent review article [4] with several historical comments on BL observations by scientists and trained professionals. The main goal of this work is to argue that the numerous puzzling observations of the ball lighting (BL) could be explained from unified viewpoint within a specific dark matter model, the so-called axion quark nugget (AQN) dark matter model [5], see also brief review [6] for a short overview of the AQN framework.
@@
Before we proceed with our explanations of the BL phenomenology within AQN framework we should, first of all, highlight the mysterious properties of the observations [1, 2, 3] which are impossible to understand if interpreted in terms of the conventional physics. In fact a complete failure to understand even the very basic features of the BL phenomenology (such as required power, or passing through a solid glass) enforced the researchers to look for possible answers to subatomic physics, well outside the conventional mechanisms considered in the past. In particular, in ref. [7] it was suggested that the magnetic monopole might be powering the BL, while in ref. [8] this idea was modified by adding an electrical charge into the system by making the dyon (magnetic monopole with non-vanishing electric charge). Our proposal is also deeply rooted into subatomic physics, but in dramatically different way, as we discuss below.
@@ 1.2 Brief overview of dark matter
We reiterate the main claim of this work slightly differently: our proposal is that the source of energy which is powering BL events is related to the annihilation events of the antimatter hidden in form of the AQNs with surrounding atoms and molecules. The fuel which is powering the BL physics is the antimatter nuggets which had been formed during the QCD transition in early Universe. It has been also shown that the dominant fraction of these antimatter nuggets mostly survive until present epoch, see brief review on formation mechanism and survival pattern in next Sect. 2.1.
This model was invented long ago to resolve fundamental problems in cosmology, not related in any way to BL phenomena (in contrast with numerous proposals which were specifically designed to explain different aspects of the BL phenomenology). Nevertheless, this AQN framework may also shed a light to another long standing problem which is the nature of the BL. We stress that all parameters for this model were fixed long ago in our previous applications to explain some mysterious and puzzling observations at galactic and solar scales, to be reviewed in Conclusion in Sect. 5.2. We keep all these parameters of the AQN model identically the same and we do not attempt to modify them to better fit the observations related to BL phenomenology.
@@ 2.1 The basics
2.2 When the AQNs hits the Earth
@@
When the AQN enters the region of the baryon density $n$ the annihilation processes start and the internal temperature increases. A typical internal temperature $T$ of the AQNs can be estimated from the condition that the radiative output must balance the flux of energy onto the nugget [33]:
@@ 2.3 AQN spallation
To conclude this overview section on AQN framework one should mention here that this model with the same set of parameters to be used in the present work may explain a number of puzzling and mysterious observations which cannot be understood as conventional astrophysical phenomena. In particular, there are many mysterious observations which occur at dramatically different scales at very different cosmological eras, which also might be related to AQN-induced phenomena. It includes BBN epoch, dark ages, as well as galactic and Solar environments. It also includes a number of puzzling and mysterious CR-like events which could be also related to AQN-induced phenomena, see concluding sections 5.1 and 5.2 for the details and references.
@@ 2. The power of radiation (total power) can be estimated as
where we use $\rm\rho(Si)=2.3g/cm^{3}$ for the estimate. The estimate (28) is very instructive as it shows a dramatic difference between two scales (21) and (28). A proper computation of the internal temperature (and corresponding photon’s energy) entering the estimate (28) is very hard technical problem, see footnote 4. However, our generic claim that the scale (28) must be dramatically smaller than the scale (21) is a solid qualitative prediction of the framework as it is entirely determined by the differences in photon’s mean free paths for two very different environments with similar x-rays energies. The estimate (28) shows the consistency of the entire framework when the scale of the BL phenomena is determined by the interaction of the $AQN_{s}$ core with environment, rather than by its internal structure. As we discussed in previous Sect. 3.4 the internal size of the $AQN_{s}$ core is dramatically smaller than scale (28).
@@ 3.6 How does BL emerge (in form of the $AQN_{s}$) after spallation?
1. We consider the numerical value for $L$ to be a very reasonable estimate. Indeed, the scale (30) corresponds to a typical size of the thunderclouds. Therefore, the $AQN_{s}$ can reach the earth’s surface after spallation in thunderclouds by loosing its huge original velocity to become BL with very low velocity at the surface where it is normally observed. A typical stoppage time is estimated as $L/v_{AQN}\sim 10^{-2}\rm s$ which is much shorter than BL’s life time $\tau$ from (26);
@@
3. The estimate (30) also shows why extremely large $B_{s}\gg 10^{15}$ have never been observed as the observed maximum for $E_{\rm max}$ is, at most, two orders of magnitude above the average value according to (1). In the AQN framework this feature is explained as follows. Very large values of $B_{s}\gg 10^{15}$ imply that the stoppage distance $L$ must be much larger in comparison with our estimate (30). Therefore, the secondary $AQN_{s}$ with very large $B_{s}$, even if they are formed, will hit the earth’s surface with very high velocities $\sim 10^{-3}c$ and get completely annihilated in deep underground regions, being unnoticed, similar to original AQN;
@@ 3.8 Concluding comments on proposal (17)
We conclude this section with the following very generic comments. The AQN model was suggested long ago to resolve some fundamental problems in cosmology, see Sect.2. The corresponding AQN parameters (including spectral properties of radiation, etc) were also worked out long ago to address many puzzles and mysteries mostly related to observed excesses of radiation at different frequencies bands at different cosmological and astrophysical scales, see Sect.5.2 for a brief review. The AQN model was not designed to address the BL physics. In our estimates in this work we use exactly the same set of parameters extracted from our cosmological studies to test bold unorthodox proposal (17) relating the BL and DM physics.
@@ 4 Frequency of appearance
where we assumed the standard halo model with the local dark matter density being $\rho_{\rm DM}\simeq 0.3\,{\rm{GeV}{cm^{-3}}}$ and canonical galactic wind $v_{\rm AQN}\simeq 220~{}\rm km~{}s^{-1}$. The number density of the AQNs is very tiny: $n_{\rm AQN}\simeq\left(\frac{\rho_{\rm DM}}{m_{p}\langle B\rangle}\right)$ as it is suppressed by factor $B^{-1}$. This should be contrasted with canonical type of WIMPs with typical mass $\sim 10^{2}\rm~{}GeV$ in comparison with AQN mass $\sim 10^{25}\rm~{}GeV$. Conventional DM detectors designed for WIMP searches are obviously useless to study the DM in form of the AQNs as there will be no any events for million of years. The Cosmic-Ray (CR) labs, on other hand, may record the AQN-induced events, and we shall comment on this with relation to recording some unusual CR -like events during the thunderstorms in Sect.5.1.

### v2

@@ Abstract
Ball lighting (BL) has been observed for centuries. There are large number of books, review articles, and original scientific papers devoted to different aspects of BL phenomenon. Yet, the basic features of this phenomenon have never been explained by known physics. The main problem is the source which could power the dynamics of the BL. We advocate an idea that the dark matter (DM) in form of the axion quark nuggets (AQN) made of standard model quarks and gluons (similar to the old idea of the Witten’s strangelets) could internally generate the required power. The AQN model was invented long ago without any relation to the BL physics. It was invented with a single motivation to explain the observed similarity $\Omega_{\rm DM}\sim\Omega_{\rm visible}$ between visible and DM components. This relation represents a very generic feature of this framework, not sensitive to any parameters of the construction. However, with the same set of parameters being fixed long ago this model is capable to address the key elements of the BL phenomenology, including the source of the energy powering the BL events. In particular, we argue that the visible size of BL, its typical life time, the frequency of appearance, etc are all consistent with suggested proposal when BL represents a profound manifestation of the DM physics represented by the AQN objects. We also argue that some of the Unidentified Aerial Phenomena (UAP) might be closely related to BL events, and therefore also represent profound manifestations of the DM physics within AQN framework. We also formulate a number of specific possible tests which can refute or unambiguously substantiate this unorthodox proposal on nature of BL and UAP.
@@ 1 Introduction
The title of this work seemingly includes two contradicting terms: the first one is “ball lighting”, which is very bright luminous object, though of unknown nature. The second term of the title is “dark matter” (DM) which is, by definition, must decouple from the radiation, as it cannot emit light. Indeed, the BL phenomenon has been known for centuries, see recent book [1] and review papers [2, 3, 4] with large number of references therein. See also recent review article [5] with several historical comments on BL observations by scientists and trained professionals. The main goal of this work is to argue that the numerous puzzling observations of the ball lighting (BL) could be explained from unified viewpoint within a specific dark matter model, the so-called axion quark nugget (AQN) dark matter model [6], see also brief review [7] for a short overview of the AQN framework.
@@
Before we proceed with our explanations of the BL phenomenology within AQN framework we should, first of all, highlight the mysterious properties of the observations [1, 2, 3, 4] which are impossible to understand if interpreted in terms of the conventional physics. In fact a complete failure to understand even the very basic features of the BL phenomenology (such as required power, or passing through a solid glass) enforced the researchers to look for possible answers to subatomic physics, well outside the conventional mechanisms considered in the past. In particular, in ref. [8] it was suggested that the magnetic monopole might be powering the BL, while in ref. [9] this idea was modified by adding an electrical charge into the system by making the dyon (magnetic monopole with non-vanishing electric charge). Our proposal is also deeply rooted into subatomic physics, but in dramatically different way, as we discuss below.
@@ 1.2 Brief overview of dark matter
We reiterate the main claim of this work slightly differently: our proposal is that the source of energy which is powering BL events is related to the annihilation events of the antimatter hidden in form of the AQNs with surrounding atoms and molecules [^1] In fact, the antimatter as a possible source of the energy powering BL has been discussed long ago [19, 20]. I thank Karl Stephan for pointing out into these two old papers.. The fuel which is powering the BL physics is the antimatter nuggets which had been formed during the QCD transition in early Universe. It has been also shown that the dominant fraction of these antimatter nuggets mostly survive until present epoch, see brief review on formation mechanism and survival pattern in next Sect. 2.1.
This model was invented long ago to resolve fundamental problems in cosmology, not related in any way to BL phenomena (in contrast with numerous proposals which were specifically designed to explain different aspects of the BL phenomenology). Nevertheless, this AQN framework may also shed a light to another long standing problem which is the nature of the BL. We stress that all parameters for this model were fixed long ago in our previous applications to explain some mysterious and puzzling observations at galactic and solar scales, to be reviewed in Conclusion in Sect. 6.3. We keep all these parameters of the AQN model identically the same and we do not attempt to modify them to better fit the observations related to BL phenomenology.
The presentation of this work is organized as follows. Next Sect. 2 represents a brief overview of the AQN construction. In our main Sect. 3 we argue that the various of observations as formulated above and collected for decades can be naturally explained within AQN framework. In Sect. 4 we estimate the BL event rate and find it is consistent with observations. In Sect. 5 we argue that some of the Unidentified Aerial Phenomena (UAP) might be closely related to BL events, and therefore also represent profound manifestations of the DM physics. In our concluding Sect. 6 we suggest a number of specific tests which can substantiate or refute our proposal on close relation between BL, UAP and dark matter physics within AQN framework.
@@ 2.1 The basics
2.2 When the AQN hits the Earth
@@
When the AQN enters the region of the baryon density $n$ the annihilation processes start and the internal temperature increases. A typical internal temperature $T$ of the AQN can be estimated from the condition that the radiative output must balance the flux of energy onto the nugget [36]:
@@ 2.3 AQN spallation
To conclude this overview section on AQN framework one should mention here that this model with the same set of parameters to be used in the present work may explain a number of puzzling and mysterious observations which cannot be understood as conventional astrophysical phenomena. In particular, there are many mysterious observations which occur at dramatically different scales at very different cosmological eras, which also might be related to AQN-induced phenomena. It includes BBN epoch, dark ages, as well as galactic and Solar environments. It also includes a number of puzzling and mysterious CR-like events which could be also related to AQN-induced phenomena, see concluding sections 6.1 and 6.3 for the details and references.
@@ 2. The power of radiation (total power) can be estimated as
where we use $\rm\rho(Si)=2.3g/cm^{3}$ for the estimate. The estimate (28) is very instructive as it shows a dramatic difference between two scales (21) and (28). A proper computation of the internal temperature (and corresponding photon’s energy) entering the estimate (28) is very hard technical problem, see footnote 5. However, our generic claim that the scale (28) must be dramatically smaller than the scale (21) is a solid qualitative prediction of the framework as it is entirely determined by the differences in photon’s mean free paths for two very different environments with similar x-rays energies. The estimate (28) shows the consistency of the entire framework when the scale of the BL phenomena is determined by the interaction of the $AQN_{s}$ core with environment, rather than by its internal structure. As we discussed in previous Sect. 3.4 the internal size of the $AQN_{s}$ core is dramatically smaller than scale (28).
@@ 3.6 How does BL emerge (in form of the $AQN_{s}$) after spallation?
1. We consider the numerical value for $L$ to be a very reasonable estimate. Indeed, the scale (30) corresponds to a typical size of the thunderclouds. Therefore, the $AQN_{s}$ can reach the earth’s surface after spallation in thunderclouds by loosing its huge original velocity to become BL with very low velocity at the surface where it is normally observed. A typical stoppage time $\tau_{s}$, when $AQN_{s}$ looses its 99% of its momentum, can be estimated as $\tau_{s}\approx\rm 2~{}s$ which is slightly shorter than BL’s life time $\tau$ from (26). At this velocity the BL becomes observable in visible frequency bands. A quantitative estimate for this stoppage time $\tau_{s}\approx\rm 2~{}s$ is given in Appendix B;
@@
3. The estimate (30) also shows why extremely large $B_{s}\gg 10^{15}$ have never been observed as BL events. Indeed, the observed maximum for $E_{\rm max}$ is, at most, two orders of magnitude above the average value according to (1). In the AQN framework this feature is explained as follows. Very large values of $B_{s}\gg 10^{15}$ imply that the stoppage distance $L$ must be much longer in comparison with our estimate (30). Therefore, the secondary $AQN_{s}$ with very large $B_{s}\gg 10^{15}$, if they are formed, will hit the earth’s surface with very high velocities and get completely annihilated only in deep underground regions, in contrast with BL which assume very low velocity near the surface. As a result, such energetic events with $B_{s}\gg 10^{15}$ are less likely to be observed in comparison with typical BL [^6] In fact, there are many recorded events, which are classified as “Pseudo-meteorites events” when a meteor-like event is observed, but no actual physical meteorite is found in the area, see 5.1 for references and details.;
@@ 3.8 Concluding comments on proposal (17)
We conclude this section with the following very generic comments. The AQN model was suggested long ago to resolve some fundamental problems in cosmology, see Sect.2. The corresponding AQN parameters (including spectral properties of radiation, etc) were also worked out long ago to address many puzzles and mysteries mostly related to observed excesses of radiation at different frequencies bands at different cosmological and astrophysical scales, see Sect.6.3 for a brief review. The AQN model was not designed to address the BL physics. In our estimates in this work we use exactly the same set of parameters extracted from our cosmological studies to test bold unorthodox proposal (17) relating the BL and DM physics.
@@ 4 Frequency of appearance
where we assumed the standard halo model with the local dark matter density being $\rho_{\rm DM}\simeq 0.3\,{\rm{GeV}{cm^{-3}}}$ and canonical galactic wind $v_{\rm AQN}\simeq 220~{}\rm km~{}s^{-1}$. The number density of the AQNs is very tiny: $n_{\rm AQN}\simeq\left(\frac{\rho_{\rm DM}}{m_{p}\langle B\rangle}\right)$ as it is suppressed by factor $B^{-1}$. This should be contrasted with canonical type of WIMPs with typical mass $\sim 10^{2}\rm~{}GeV$ in comparison with AQN mass $\sim 10^{25}\rm~{}GeV$. Conventional DM detectors designed for WIMP searches are obviously useless to study the DM in form of the AQNs as there will be no any events for million of years. The Cosmic-Ray (CR) labs, on other hand, may record the AQN-induced events, and we shall comment on this with relation to recording some unusual CR -like events during the thunderstorms in Sect.6.1.
@@
5 On Possible relation between BL and other Unidentified Aerial Phenomena (UAP)
In this section we would like to present several arguments suggesting that some of the Unidentified Aerial Phenomena (UAP) might be also related to the DM physics. For a comprehensive review on UAP observations see recent review [47].
By obvious reasons this section is much more speculative in comparison with all our previous sections because of scarcely real physical measurements of UAP events in comparison with sufficiently long list of numerical estimates for BL events as reviewed in Sect. 1.1. In fact, items (ix), (x) and (xi) from Sect. 1.1 can be treated as the BL-features recorded by modern physics instruments. This should be contrasted with UAP-related phenomena when very few observations could be considered as well documented and properly recorded events. In fact, scientific studies of the UAP phenomena started very recently, see e.g. papers [48, 49, 50]. Nevertheless, we opted to include descriptions of several such observations into this work because we believe that some of the UAP observations might be closely related to BL physics. Therefore, some UAP events may also represent the manifestations of the DM physics as a consequence of our proposal (17).
Indeed, there is one unique but crucial common element relating BL phenomena and UAP: in both cases no any material objects remain in the active area where events took places. This is in spite of the fact that many secondary accompanying effects had been observed and even measured in physics laboratories. The secondary effects manifested in many different ways: in form of the visible light, X rays, shock acoustic waves (see Sect. 5.2), heat melting the glass or soil (see Sect.5.1). This is only possible when the source of the energy for both cases is represented by annihilation processes of the antimatter with surrounding matter materials, which is precisely the key element of this proposal.
As a result of these similarities we propose that some of the UAP events may be also considered as profound manifestations of the same DM physics similar to our proposal (17) relating BL and DM physics. Therefore, the main purpose of this section is to argue that many UAP events are in fact very similar to the BL phenomena discussed in Sect. 3. The only difference is that the UAP events are characterized by dramatically different energy scales (larger baryon charge, and consequently much higher velocity at the moment of observation) in comparison with BL events with low velocities and typical energies (18).
5.1 Pseudo-meteorites events as the AQN-induced events
There are many reports in the literature describing the meteor-like events which (after detail studies by professionals) turned out to be not the meteorites. The corresponding events are classified as “Pseudo-meteorites events”, see [51] with large number of references and details. In this work we want to focus on a single event which took place in US town of Elma (Washington State) on July 15, 2003 as described in [51]. Many other similar events discussed in [51] shall not be mentioned here, and we think they likely have the same nature, and are originated from the same physics.
The key elements of the Elma event are [51]: 1.after the event (fireball) the glassy black rocks had been found; 2. the rocks were very hot (one of the witnesses had burned a thumb and a finger by collecting the rocks); 3. analysis of the discovered stones by the University of Washington concluded that the glassy rocks were not meteorites, see pictures in [51]; 4. security cameras did not record any suspicious images; 5. cloud of dust went up in the active area.
Our original comment is as follows. The Elma event is very similar to BL events discussed in Sect. 3 with the only difference is that the $AQN_{s}$ must be much larger in size $B\gg 10^{15}$ in comparison with a typical BL event. In this case, as we discussed in item 3 in Sect. 3.6 the $AQN_{s}$ cannot efficiently slowdown, and it is likely to hit the earth’s surface with a high speed. As a result, the $AQN_{s}$ gets annihilated in deep underground regions (not in air, which is typical for BL events) by heating the surrounding rocks along its path. In many respects the heating of the surrounding material is similar to phenomena of the BL passing through the glass window when the internal $AQN_{s}$ temperatures could reach enormous values due to much higher density of the soil in comparison with air, see our estimates in Sect. 3.5. It explains the hot glassy rocks found in the area. Furthermore, for the $AQN_{s}$ moving with a high speed the radiation in visible frequency bands is suppressed (similar to ELFO event discussed in next Sect.5.2), in contrast with slow moving BL discussed in Sect. 3.3. It explains the absence of any images by security camera. This identification of the “Pseudo-meteorites events” with $AQN_{s}$ events is consistent with all observed puzzling observations listed in previous paragraph. Therefore, we speculate that the “Pseudo-meteorites events” along with BL events represent the right hand side of the same identification (17) and both types of events represent profound manifestations of the same DM physics.
5.2 Sky-quakes as the AQN events
The so-called skyquakes have been known for centuries, similar to BL events. They manifested in the form of sound and infrasound without leaving any physical material objects in the area in form of true meteorities. Their nature remains unknown in spite of the long history of observations, with records going back over 200 years. It has been speculated in [31] that skyquakes could be a manifestation of the dark matter AQN traveling in the atmosphere.
Unfortunately, it is next to impossible to extract any useful quantitative information from the numerous but random and unsystematic records on sky-quakes collected for centuries (very similar to BL events). Luckily, one such event which occurred on July 31-st 2008 was properly recorded by the dedicated Elginfield Infrasound Array (ELFO) near London, Ontario, Canada, see [31] for references and details. The infrasound detection was accompanied by non-observation of any meteors by an all-sky camera network, ruling out a conventional meteor source. In addition, no any meteorites had been found in the active area, similar to Elma event mentioned above in Sect. 5.1 and classified as “Pseudo-meteorites events”. Anthropogenic sources such as operations at the nearby Bruce Nuclear Power Plant or the Goderich salt mine were also eliminated; a local airport radar reported no aircraft in the area at the time. In addition to infrasound, impulses were also observed seismically (few moments after the event) as ground-coupled acoustic waves around Southwestern Ontario and Northern Michigan. This event was treated as AQN-induced event as discussed in detail in [31] where it was argued that the energetics, infrasound-frequency properties, and other characteristics of the event are consistent with observed ELFO event.
Our original comment here is as follows. As argued in [31] the skyquakes could be also AQN induced events. However, the scale of skyquakes are dramatically different from BL events and “Pseudo-meteorites events” mentioned above. In fact, it was estimated in [31] that the the baryon charge of ELFO event is around $B\approx 10^{27}$ which is two orders of magnitude above a typical AQN size [^8] Such powerful events are much more rare ones than typical AQN events with rate given by (8) because the size distribution $f(B)\propto B^{-2}$ and many orders of magnitude above a typical $AQN_{s}$ scale. In different words, skyquakes are the manifestations of the original AQN, not secondary $AQN_{s}$ with dramatically smaller sizes and masses (and consequently released energies). However, all types of events discussed above (BL, Pseudo-meteorites, and skyquakes) can be considered as different manifestations of the same AQN-induced events when the source of the energy is the same, and it is hidden in form of the antimatter nuggets which had been produced during the QCD epoch in early Universe.
5.3 Unidentified Anomalous Vehicles (UAV) as manifestation of the AQN events
This is the most speculative subsection of this work as no quantitative physical measurements exist of the so-called Unidentified Anomalous Vehicles (UAV) which have been observed globally, see recent review [47] and scientific publication [52] with specific numerical estimates for the velocity and acceleration of UAV which seem to be inconsistent with laws of physics if interpreted as aircraft.
We interpret the radar information of the UAV image from USS Nimitz nuclear aircraft carrier (14 November 2004), see Fig. 5 in [52], as an image of a highly ionized cloud (cylindrical form) produced by passing AQN. In what follows we support this interpretation by estimation of the plasma frequency $\omega_{p}$ of this ionized cloud. In this case the pulse being sent from radar will be reflected by this highly ionized cloud and recorded by USS Nimitz. The effect of reflection from ionized cloud is similar to reflection of the AM radio bands from the Earth’s ionosphere. In contrast with ionosphere however where $\omega_{p}\approx 2\cdot 10^{7}s^{-1}$ (MHz range), the plasma frequency in highly ionized cloud produced by passing AQN could easily assume GHz range. As a result, the AQN-induced cloud could be misinterpreted as UAV image. In what follows we support our proposal by providing an order of magnitude estimates for $\omega_{p}$ and the size of the highly ionized cloud produced by passing AQN.
Indeed, our order of magnitude estimates as given in Appendix C suggest that the AQN propagating at high altitudes will be ionizing the surrounding area with very high efficiency. We estimated the plasma frequency for this ionized region to be on the level $\omega_{p}\approx 3\cdot 10^{9}\rm s^{-1}$ or even much higher, according to (52). This implies that the pulses with GHz frequency bands from radar can be reflected by the ionized cloud999 One should emphasize that an ionized cloud produced by passing AQN is the subject of atmospheric currents and can easily move and fluctuate. The AQN itself entering the atmosphere with very high speed $\sim 10^{-3}c$ could be at a very different location at the instant when the radar records the image of the ionized cloud., which can be misinterpreted as UAV. Furthermore, a typical size of the cloud is determined by the mean free path of the 20 keV photons emitted by very hot AQN. The corresponding mean free path is estimated in (50) and it assumes the numerical value around $L_{\gamma}\approx\rm 15~{}m$. This scale is similar to the size of UAV image mentioned in ref. [52]. Finally, the acceleration and the velocity of the AQN propagating in atmosphere have been estimated in (44) and (47) and consistent with the values presented in ref. [52]. The estimation (47) shows that the acceleration could easily reach enormous magnitudes $\sim(10^{2}-10^{4})g$ being inconsistent with modern technology if interpreted as UAV according to ref. [52]. Of course, such acceleration is perfectly consistent with physics laws if interpreted as the AQN object which enters the earth’s atmosphere with galactic velocity (around $10^{-3}c$) and slows down as a result of collisions with atoms and molecules in atmosphere as described in Appendix B.
We conclude this section by mentioning that UAV along with BL, Pseudo-meteorites, and skyquakes as discussed above represent different manifestations of the same AQN-induced events when the source of the energy is hidden in the antimatter nuggets which had been formed during the QCD epoch in early Universe. All these events are very rare ones. Their rareness is explained by very tiny DM flux formulated in terms of the well-constrained DM density $\rho_{\rm DM}$ as given by (33).
It is obvious that we can not make any progress in understanding of the UAP events mentioned above due to the lack of real data and proper physics measurements, which makes this section even more speculative than BL studies with sufficiently long list of numerical estimates as reviewed in Sect. 1.1. Fortunately, the situation is about to change, and a number of projects and proposals, see e.g. [48, 49, 50] have been put forward to fill this gap. The UAP events in general and UAV events in particular can be and should be systematically studied. We propose in Sect. 6.2 several specific tests on how to substantiate or refute our proposal formulated above on identification some of the UAP events with AQN-induced phenomena.